OpenStudy (anonymous):

Please VERIFY (cscx-1)(sinx+1)=cosxcotx

OpenStudy (anonymous):

I want to work the left side... (cscx-1)(sinx+1)= csc(x)sin(x)+csc(x)-sin(x)-1= 1+csc(x)-sin(x)-1= csc(x)-sin(x)= 1 / sin(x) - sin(x)= 1 / sin(x) - sin^2(x) / sin(x)= 1 - sin^2(x) / sin(x)= cos^2(x) / sin(x)= [ cos(x) / sin(x) ] * cos(x) = [ cot(x) ] * cos(x) = both sides are equal.

OpenStudy (anonymous):

Wow that makes perfect sense. Thanks!

OpenStudy (anonymous):

You welcome !